OpenStudy (anonymous):

The cube shown has a volume of 125cm3.What is the area? A.I.M.s sample test

OpenStudy (anonymous):

A: 15cm2 B: 20cm2 C:25 cm2 D:30 cm 2 Thanks!

OpenStudy (anonymous):

Volume of cube = (side)^3 Volume of cube = 125 cm^3 (side)^3 = 125 (side)^3 = (5)^3 side = 5 cm Area = (side)^2 Area = (5)^2 Area = 25 cm^2